Two forces , one four times as large as the other pull in the same direction on a 10kg mass and impart to it an acceleration of 2.5 m/s ^ 2 ?. If the smaller force is removed , what is the acceleration of the mass?
1
Expert's answer
2020-12-04T11:16:42-0500

Let the smaller force be "F" and the bigger one be "4F". Since they are pulling in the same direction, the net force is "F+4F = 5F". Then, according to the Newton's second law:


"5F = ma_0"

where "m = 10kg" is the mass of the object and "a_0 = 2.5m\/s" is its acceleration. Thus, the smaller force is:


"F = \dfrac{10\cdot 2.5}{5} = 5N"

If we remove the smaller force, only the "4F" will be left. Then, using the Newton's second law again, obtain:


"4F = ma"

where "a" is the unknown acceleration with only bigger force left. Thus, obtain:


"a = \dfrac{4F}{m} = \dfrac{4\cdot 5}{10} = 2m/s^2"

Answer. "2\space m/s^2".